OpenStudy (anonymous):

Find the solution to the equations. 3x - y = -4 x + y = 0

OpenStudy (anonymous):

you can use elimination with these equations so you add both 3x-y=-4 +x +y=0 4x=-4 divide by 4 x=-1 then substitute x into either equation i will show the 2nd -1+y=0 y=1

OpenStudy (austinl):

Easy peasy! We have, 3x - y = -4 x + y = 0 Solve the second equation for y, x + y = 0 y = -x Plug that in the first equation, 3x - (-x) = -4 3x+x=-4 4x=-4 x=-1 Then you can plug that into the other equation, x + y = 0 (-1) + y = 0 y = 1